OpenStudy (anonymous):

Solve for w. 5w + 9z = 2z + 3w

OpenStudy (texaschic101):

5w + 9z = 2z + 3w --- subtract 3w from both sides 5w - 3w + 9z = 2z --- subtract 9z from both sides 5w - 3w = 2z - 9z --- combine like terms 2w = - 7z --- divide by 2 on both sides w = -7/2z

OpenStudy (texaschic101):

any questions at all ?

OpenStudy (olyalhic):

5w+9z=2z+3w Add -3w to both sides. 5w+9z+−3w=3w+2z+−3w 2w+9z=2z Add -9z to both sides. 2w+9z+−9z=2z+−9z 2w=−7z Divide both sides by 2. \[ \frac{ 2w }{ 2 }= \frac{ -7z }{ 2 }\] Answer w=−72z

OpenStudy (olyalhic):

* -\[-\frac{ 7 }{ 2 }z\]